# Post-hoc Study of Climate Microtargeting on Social Media Ads with LLMs: Thematic Insights and Fairness Evaluation

### v3

@@ 6.1 Fairness Analysis on Gender Prediction
Equal Opportunity (True Positive Rate) focuses on the True Positive Rates (TPR) across gender groups, measuring the model’s ability to correctly identify positive instances within each group. We have Female TPR: $\mathbf{0.95}$ and Male TPR: $\mathbf{0.85}$. The TPR for females is higher by $0.10$, indicating that model may be more effective at correctly identifying females than males.
@@
Table 9: Equal Opportunity (TPR) and Predictive Equality (FPR) for Age Groups.
@@ 6.2 Fairness Analysis on Age Prediction
Table 9 shows that the TPR for the senior group is notably lower, suggesting the model is less effective at correctly identifying individuals in this age group. The senior group has an FPR of 00, indicating no false positives, while the early working group has the highest FPR among the groups.
@@ 6.3 Statistical Significance Tests
While a 0.10 difference in TPR may seem modest, prior fairness literature Hardt et al. (2016) recognizes it as meaningful in sensitive domains. We conduct two-proportion Z-tests to assess whether the observed differences in TPRs across demographic groups are statistically significant. For gender prediction, the difference in TPR between females ($0.95$) and males ($0.85$) resulted in a p-value of $0.086$ with a $95\%$ confidence interval of $[-0.018,0.230]$. This result indicates that the gap is not statistically significant at the conventional $0.05$ level. However, in the age group analysis, the TPR difference between young adults ($0.88$) and seniors ($0.33$) yields a p-value of $0.004$ with a $95\%$ confidence interval of $[0.138,0.796]$, indicating a statistically significant disparity. This validates our concern about fairness in age-related predictions, particularly for underrepresented senior groups.
@@ 6.4 Sensitivity Analysis
Fig 7(a) in App. A.6 shows that bias persists after conditioning on ad spend. Low-budget ads: $\Delta\approx 0.82$ ($95\%$ CI 0.680.680.680.680.960.960.960.96), High-budget ads: $\Delta\approx 0.76$ ($95\%$ CI 0.550.550.550.550.930.930.930.93). Overlapping CIs indicate similar magnitude and ad-budget is not driving the disparity. Topic clusters tell the same story (Fig 7(b) in App. A.6). Four of five clusters show sizable positive gaps (0.300.300.300.301.001.001.001.00). The one cluster with only female-authored ads (topic 2) cannot yield a gap, but its omission does not overturn the pattern. Accuracy gaps are modest and hover around zero. Where CIs cross zero, classification accuracy is comparable between genders; the selection-rate bias remains the dominant concern.
[…]
@@ 6.5 Robustness Analysis
To provide a more in-depth evaluation of the quality and robustness of the data, we plot learning curve containing macro-F1 and demographic-parity against sample size (20202020227227227227 ads). Learning-curve plot in Fig. 8 in App. A.7 shows that macro-F1 and fairness gaps plateau well before 227 examples—evidence that the analysis is not variance-limited.
@@
Figures 8(c) and 8(d) in App. A.7 show the macro-F1 and age-parity gap at 20-ad increments. Macro-F1 rises to $0.70$ by $n\approx 80$ and then plateaus; the $95\%$ confidence band narrows to ±$0.8$ pp. The parity gap likewise converges to $\approx 0.57$ with error $<0.02$. Thus, our age-bias findings are not driven by sample variance—adding more data in the current distribution neither boosts accuracy nor mitigates disparity.
@@ 7 Conclusion
Our work advances the discourse on climate microtargeting by demonstrating the utility of LLMs in accurately detecting and explaining targeted messaging strategies on social media. We show our analysis on climate campaigns dataset, but our approach can easily be adapted in any dataset. It is designed to be scalable without any modifications. We provide a broader theme and various aspects under each theme based on the explanations from LLMs, which we hope will be an important contribution to the CSS community. Besides, the fairness analysis conducted in our study underscores the importance of evaluating and addressing the biases. Disparities in prediction accuracy and error rates, highlight the need for more inclusive and equitable targeting methods.
@@ 8 Limitations
We have small sample size due to restrictive filtering criteria which might result in underrepresented groups, particularly late working and senior demographics. The rationale for using this specific dataset is that it is publicly available and, to the best of our knowledge, it is the only publicly available dataset focused on targeted climate campaigns.
@@ 9 Ethics Statement
To the best of our knowledge, we did not violate any ethical code while conducting the research work described in this paper. We report the technical details for the reproducibility of the results. In this paper, we did not introduce any new dataset; instead, we experimented using existing dataset that are adequately cited. The author’s personal views are not represented in any qualitative results we report, as it is solely an outcome derived from machine learning models.
@@ A.2 Demographic Misclassifications
A.2.1 Misclassification - Senior Instances
In Table 11, the fourth instance shows a misclassification where the model predicts it as a late working (45-64 years) age group. The ad mentions ‘working families’, which is a term commonly associated with individuals in the Late Working age group who are actively engaged in the workforce and supporting families. The content revolves around a political campaign emphasizing the need for change and active participation, themes often associated with the $45-64$ age demographic who are typically more politically active and influential.
@@ A.2.1 Misclassification - Senior Instances
Get clean water and pollution-free electricity to all of America. Young Targets environmentally conscious young adults concerned about sustainability and future impact of clean water and energy.
@@ A.5 Unsupervised Baselines
Generated topics from prediction explanation of gender and age using LDA and BERTopic is shown in Table 13. Table 12 shows the human evaluation results on GPT-4o assignment, LDA, and BERTopic. Prompt template for LLM assignment is shown in Fig. 6.
@@
Topic 4: women, ad, environmental, making, issues, future, likely, mothers, childrens, concerned

### v4

@@ 6.1 Fairness Analysis on Gender Prediction
Equal Opportunity (True Positive Rate) focuses on the True Positive Rates (TPR) across gender groups, measuring the model’s ability to correctly identify positive instances within each group. We have Female TPR: $\mathbf{0.95}$ and Male TPR: $\mathbf{0.85}$. The TPR for females is higher by $0.10$, indicating that the model may be more effective at correctly identifying females than males.
@@
Table 9: TPR and FPR for Age Groups.
@@ 6.2 Fairness Analysis on Age Prediction
Table 9 shows that the TPR for the senior group is notably lower, suggesting the model is less effective at correctly identifying individuals in this age group. The senior group has an FPR of 0, indicating no false positives, while the early working group has the highest FPR among the groups.
@@ 6.3 Statistical Significance Tests
While $0.10$ difference in TPR may seem modest, prior fairness literature Hardt et al. (2016) recognizes it as meaningful in sensitive domains. We conduct two-proportion Z-tests to assess whether the observed differences in TPRs across demographic groups are statistically significant. For gender prediction, the difference in TPR between females ($0.95$) and males ($0.85$) resulted in a p-value of $0.086$ with a $95\%$ confidence interval of [0.018,0.230][-0.018,0.230]. This result indicates that the gap is not statistically significant at the conventional $0.05$ level. However, in the age group analysis, the TPR difference between young adults ($0.88$) and seniors ($0.33$) yields a p-value of $0.004$ with a $95\%$ confidence interval of $[0.138,0.796]$, indicating a statistically significant disparity. This validates our concern about fairness in age-related predictions, particularly for underrepresented senior groups.
@@ 6.4 Sensitivity Analysis
Fig 7(a) in App. A.6 shows that bias persists after conditioning on ad spend. Low-budget ads: $\Delta\approx 0.82$ ($95\%$ CI 0.680.680.960.96), High-budget ads: $\Delta\approx 0.76$ ($95\%$ CI 0.550.550.930.93). Overlapping CIs indicate similar magnitude, and ad-budget is not driving the disparity. Topic clusters tell the same story (Fig. 7(b) in App. A.6). Four of five clusters show sizable positive gaps (0.300.301.001.00). The one cluster with only female-authored ads (topic 2) cannot yield a gap, but its omission does not overturn the pattern. Accuracy gaps are modest and hover around zero. Where CIs cross zero, classification accuracy is comparable between genders; the selection-rate bias remains the dominant concern.
Fig 7(c) in App. A.6 illustrates that the gap is not a low-budget artifact: whether an ad costs $\mathdollar 5$ or $\mathdollar 500$, the model is still far more likely to assign it to some age groups than others. Four of five topics show large positive parity gaps—even where sample sizes differ. In Fig. 7(d) in App. A.6, Topic 1 does not contradict the bias claim as it simply contains ads that all target (and therefore are predicted for) one single age group. Stratified sensitivity checks confirm that age-group disparities persist across both advertising-budget tiers ($\Delta\approx 0.47$$1.0$) and four of five latent topic clusters, indicating that neither spend nor topic accounts for the bias.
@@ 6.5 Robustness Analysis
To provide a more in-depth evaluation of the quality and robustness of the data, we plot a learning curve containing macro-F1 and demographic-parity against sample size (2020227227 ads). Learning-curve plot in Fig. 8 in App. A.7 shows that macro-F1 and fairness gaps plateau well before 227 examples—evidence that the analysis is not variance-limited.
@@
Figures 8(c) and 8(d) in App. A.7 show the macro-F1 and age-parity gap at 20-ad increments. Macro-F1 rises to $0.70$ by $n\approx 80$ and then plateaus; the $95\%$ confidence band narrows to ±0.8±0.8 pp. The parity gap likewise converges to $\approx 0.57$ with error <0.02<0.02. Thus, our age-bias findings are not driven by sample variance—adding more data in the current distribution neither boosts accuracy nor mitigates disparity.
@@ 7 Conclusion
This study shows that LLMs can function as practical, third-party auditors of microtargeted climate ads—identifying whom messages are meant to reach, explaining why, and surfacing equity concerns that are otherwise opaque. On our climate-ad corpus, the LLMs accurately infer demographic targets overall ($88.55\%$ accuracy), with especially strong performance for gender ($94.92\%$ for female-targeted ads; $85.10\%$ for male-targeted ads) and more variable results across age groups. Beyond prediction, model-generated explanations enable a compact, reusable taxonomy of themes used to engage different audiences, providing interpretable signals that researchers and practitioners can inspect and act on. Besides, the fairness analysis conducted in our study underscores the importance of evaluating and addressing the biases. Disparities in prediction accuracy and error rates highlight the need for more inclusive and equitable targeting methods. Although we show our analysis on the climate campaigns dataset, our approach can easily be adapted to any dataset. It is designed to be scalable without any modifications. Code and dataset are available [^9] https://github.com/tunazislam/llms-posthoc-climate.
@@ 8 Limitations
We have a small sample size due to restrictive filtering criteria, which might result in underrepresented groups, particularly late working and senior demographics. The rationale for using this specific dataset is that it is publicly available and, to the best of our knowledge, it is the only publicly available dataset focused on targeted climate campaigns.
@@ 9 Ethics Statement
To the best of our knowledge, we did not violate any ethical code while conducting the research work described in this paper. We report the technical details for the reproducibility of the results. In this paper, we did not introduce any new dataset; instead, we experimented using an existing dataset that is adequately cited. The data do not contain personally identifiable information and report engagement patterns at an aggregate level. The author’s personal views are not represented in any qualitative results we report, as it is solely an outcome derived from machine learning or AI models.
10 Acknowledgments
We are grateful to the anonymous reviewers for their insightful comments. This work was partially funded by the Purdue Graduate School Summer Research Grant (to TI) and NSF CAREER award IIS-2048001.
@@ A.1 Prompting
A.2 Demographic Misclassifications
A.2.1 Misclassification - Senior Instances
In Table 11, the fourth instance shows a misclassification where the model predicts it as a late working (45-64 years) age group. The ad mentions ‘working families’, which is a term commonly associated with individuals in the Late Working age group who are actively engaged in the workforce and supporting families. The content revolves around a political campaign emphasizing the need for change and active participation, themes often associated with the 456445-64 age demographic who are typically more politically active and influential.
@@ A.2.1 Misclassification - Senior Instances
Get clean water and pollution-free electricity to all of America. Young Targets environmentally conscious young adults concerned about sustainability and the future impact of clean water and energy.
@@ A.5 Unsupervised Baselines
Generated topics from the prediction explanation of gender and age using LDA and BERTopic is shown in Table 13. Table 12 shows the human evaluation results on GPT-4o assignment, LDA, and BERTopic. Prompt template for LLM assignment is shown in Fig. 6.
@@
Topic 4: women, ad, environmental, making, issues, future, likely, mothers, children, concerned